i need to indentify the slope and y intercept of the line with this equation 3x-8y=-24

Answer by rfer(10417)

y=mx+b
3x-8y=-24
-8y=-3x-24
y=3/8x+3
m=slope=3/8
b=y int=3
(0,3)